pts) boat of mass m is vo o on a lake at x0 0 m. 2
Answer #1

2.a)Force, F(v)=-av1/2

We know that force= mass * acceleration.

Let the acceleration be A.

Therefore, F(v)=mA=- av1/2

                      mA=- av1/2

                        A=(-av^1/2)/m

Acceleration can be expressed as

                          \mathbf{A=v\frac{\mathrm{d} v}{\mathrm{d} x}}

                       \mathbf{v\frac{\mathrm{d} v}{\mathrm{d} x}=-av^{1/2}/m}

                      \mathbf{v^{1/2}dv=-\frac{adx}{m}}

     Integrating both side,

                    \mathbf{2/3v^{3/2 }-2/3v_{o}^{3/2}=-\frac{ax}{m}}

                  \mathbf{x=-\frac{m}{a}(2/3v^{3/2 }-2/3v_{o}^{3/2})}

This is the equation of motion that we get.

2.b)

    Acceleration is also defined as

                                                           \mathbf{A=\frac{\mathrm{d} v}{\mathrm{d} t}}

                                                            \mathbf{\frac{\mathrm{d} v}{\mathrm{d} t}=-av^{1/2}/m}

                                                            \mathbf{\frac{dv}{v^{1/2}}=\frac{-a}{m}dt}

                                              Integrating both side,

                                          \mathbf{2v^{1/2}-2v_{o}^{1/2}=\frac{-at}{m}}

                                          \mathbf{v =(v_{o}^{1/2}-\frac{at}{2m}})^{2}